A jet travels 6096mi against a jetstream in 8 hours and 7056mi with the jetstream in the same amount of time. What is the rate of the jet in still air and what is the rate of the jetstream?
Answer by mananth(16946)


Let Plane speed= x mph
let wind speed= y mph
with wind , speed= (x+y)
against wind, speed (x-y)
d= 7056 with wind d/s =t

7056 / (x+y) = 8
divide by 8

882 / (x+y) = 1
(x+y) = 882 ............1

d= = 6096 against wind
6096/(x-y) = 8
divide by 8

762 /(x-y) = 1
x - y = 762 .............2
add up (1) & (2)
2 x = 1644
/ 2
x= 822 mph speed of Plane in still air

plug value of x in (1)
we get y= 60 mph speed of wind